### v1

# Discovering Properties of Inflectional Morphology in Neural Emergent Communication

Abstract
Emergent communication (EmCom) with deep neural network-based agents promises to yield insights into the nature of human language, but remains focused primarily on a few subfield-specific goals and metrics that prioritize communication schemes which represent attributes with unique characters one-to-one and compose them syntactically. We thus reinterpret a common EmCom setting, the attribute-value reconstruction game, by imposing a small-vocabulary constraint to simulate double articulation, and formulating a novel setting analogous to naturalistic inflectional morphology (enabling meaningful comparison to natural language communication schemes). We develop new metrics and explore variations of this game motivated by real properties of inflectional morphology: concatenativity and fusionality. Through our experiments, we discover that simulated phonological constraints encourage concatenative morphology, and emergent languages replicate the tendency of natural languages to fuse grammatical attributes.
Discovering Properties of Inflectional Morphology in Neural Emergent Communication
Miles Gilberti Shane Storks Huteng Dai University of Michigan {milgil,sstorks,huteng}@umich.edu
1 Introduction
Emergent communication (EmCom) with deep neural agents studies communication protocols developed between two or more agents engaged in a task requiring successful communication. Analyses of learned protocols in EmCom typically search for characteristics deemed crucial in natural language, prominently compositionality Chaabouni et al. (2020). Such experiments promise to shed light on the pressures underlying human-like language evolution.
However, as interpreting emergent communication schemes and measuring compositionality are challenging, inhibitive assumptions are often made. First, most EmCom works afford agents an over-complete vocabulary of characters, making it possible for agents to describe each possible attribute value with a unique character one-to-one, thus leading to treatment of generated characters exclusively as words composed only syntactically. In turn, common evaluation metrics developed for this large-vocabulary setting are overly simplified, favoring protocols that represent attribute values with consistent characters concatenated in a consistent ordering Chaabouni et al. (2020); Resnick et al. (2020); Ueda et al. (2023). This disregards the possibility of many morphological phenomena attested in natural languages, such as nonconcatenative morphology (e.g., root-pattern morphology common in Semitic languages), and fusion of grammatical attributes in morphemes (e.g., the Spanish verb suffix -amos simultaneously communicates first person, plurality, and present tense).
In this work, we reinterpret EmCom as emergent morphology, addressing these shortcomings with two key design choices. First, we target humanlike double articulation Hockett (1960), the property of natural language that meaningful components are composed of combinations of individually meaningless components, by enforcing a small character inventory for messages, thus necessitating the combination of characters into meaningful morpheme-like symbols to express attribute values. Second, as shown in Figure 1, we introduce a novel configuration of attribute-value reconstruction inspired by inflectional morphology, where a high-cardinality attribute (analogous to roots) is paired with comparatively lower-cardinality attributes (analogous to grammatical information, e.g., tense and person), encouraging richer morphological phenomena. We review EmCom evaluation metrics for various properties of communication schemes, proposing novel metrics for the underexplored properties of concatenativity and fusionality. To validate and establish expected ranges for these metrics, we then analyze a representative group of artificial double articulatory communication schemes and inflection-based natural language communication schemes.
Lastly, we leverage this reimagined problem to explore several novel, linguistically compelling research questions. We implement a phonology-inspired constraint in communication agents, using concatenativity metrics to reveal ease of articulation as a pressure toward concatenative languages. We then compare the topographic similarity of emergent and natural languages, finding significant room for improvement in emergent language to reflect natural degrees of compositionality. Finally, we analyze the presence of fusion in emergent and natural languages for inflection, finding that emergent languages most often fuse grammatical attributes rather than roots, mirroring natural language. Altogether, this work lays a strong foundation for future research into the evolution of morphology.
2 Related Work
We adopt the term “emergent morphology” introduced by Archangeli and Pulleyblank (2016) within their Emergent Grammar framework, which hypothesizes how morphophonological structures arise from constraint interactions in natural languages. This is one of few attempts to model the emergence of aspects of morphological systems, including work by Elsner et al. (2020), who simulate how phonological processes can lead agglutinative systems to become fusional over time. Also related are formal and evolutionary analyses of double articulation conducted by Nowak and Krakauer (1999) and Zuidema and de Boer (2009). Unlike these works, we emphasize computational modeling of artificial communication systems, focusing specifically on how minimal meaningless discrete units combine systematically into meaningful morphological structures like inflection.
There exist some works that address double articulation explicitly via selection of small inventory sizes Ueda et al. (2023); Ueda and Washio (2021); Ueda and Taniguchi (2024), but the analysis focuses on words rather than morphemes, and only considers perfectly concatenative methods of combining morphemes. Work utilizing bitstring messages in EmCom also achieve this Resnick et al. (2020); Gupta et al. (2020), albeit not in a linguistically realistic way.
Other works use large-vocabulary EmCom to investigate various phenomena relevant to natural language morphology. For example, Chaabouni et al. (2019) use a length pressure to rederive Zipf’s Law of Abbreviation, a statistical property of natural language. Lian et al. (2023) investigate the specific morphosyntactic problem of the case-marking versus word-order tradeoff with pre-learned languages as a starting point, with Lian et al. (2024) extending this inquiry to a group communication setting. Conklin and Smith (2023) discover naturalistic variability in emergent languages, complicating traditional measures of compositionality. While Galke et al. (2022) argue the importance of learnability, generalization, and group size pressures in EmCom, Galke and Raviv (2024) further review the impact of learnability and production effort among other psycho-sociolinguistic factors in EmCom. Meanwhile, Dekker (2024) models the evolution of specific languages’ inflectional morphology by pretraining agents with those languages, whereas we focus on the emergence of inflectional morphology systems from scratch.
Lastly, there is extensive literature on automated morphological segmentation, and segmenters for nonconcatenative morphology are particularly relevant Fullwood (2018). While we only explored concatenative approaches, as emergent morphology develops, we expect these techniques to play a crucial role in understanding emergent languages.
3 Toward Emergent Morphology
Prior work in EmCom has largely focused on the Lewis Signaling Game Kottur et al. (2017); Chaabouni et al. (2020); Ueda et al. (2023); Lewis (1979), in which a sender agent observes some state and communicates via a message to a receiver agent, which must take an action based on this information. Success in the game relies on accurate communication of the relevant features of the environment. Morphology has been relatively understudied in EmCom due to a focus on individual symbols representing atomic units of meaning Boldt and Mortensen (2024b).
3.1 Attribute-Value Reconstruction Game
The Attrbute-Value Reconstruction Game (henceforth Attr-Val) has been extensively studied by EmCom researchers Kottur et al. (2017); Chaabouni et al. (2020). This class of games models the state as an -tuple of values where each value belongs to some attribute , defined as a set of possible values for the attribute. That is, the set of states is .
Formally, an instance of Attr-Val can be defined as such: we have a state , a sender , and a listener . is the message space, which consists of messages of up to length characters, all of which must belong to a fixed character vocabulary . This instance is considered successful if . Following prior works, we use EGG Kharitonov et al. (2019) to implement sender and receiver agents as single layer gated recurrent unit (GRU; Cho et al., 2014) models. More details about agent implementation and training are provided in Appendix F.
3.2 Attr-Val Configurations as Inflectional Morphology
The attribute value sets tend to have the same cardinality in prior work Kottur et al. (2017); Chaabouni et al. (2020); Ueda et al. (2023). So, as a conceptual framework, this setting has commonly been compared to communicating about objects with various properties e.g., a red square or blue triangle.
We claim that Attr-Val games also represent a rich analogy for inflectional morphology. The meaning of an inflected root in natural language consists of the root as well as additional attributes like tense and person for verbs, or number and gender for nouns. Like Attr-Val, these attributes take on a finite set of values. While the number of possible roots is unbounded in natural language, Attr-Val can simulate this by having one attribute set (representing roots) be much larger than others. This paradigm provides a key advantage: natural languages solve an analogous and similarly challenging problem to this, enabling direct comparison with learned communication schemes.
We consider two configurations for , both with three attribute sets: a default setting with 16 possible values per attribute (comparable to prior work), and a novel inflection setting with attribute set cardinalities of 42, 2, and 3, respectively symbolizing roots, tenses, and person.111Note that we also conducted preliminary experiments controlling for total number of combinations (683 2 3), but models universally failed to converge, leading us to conclude that symbolic complexity Zhang (2025) is a more meaningful and necessary quantity to control for. We further discuss this issue and justify our choices of values in Appendix I. Both settings have a comparable total number of attribute values (48 and 47). To necessitate double articulation, i.e., the composition of meaningful symbols from meaningless characters to represent attributes, we enforce a consistent character vocabulary size and message length , smaller than the cardinality of most attribute sets. Importantly, one attribute in the inflection setting has a cardinality much greater than while others have fewer possible values than the vocabulary size, possibly incentivizing learning of more morphologically diverse communication schemes than the default setting.
4 Interpreting Communication Schemes
To investigate the emergence of double articulation in EmCom, we next introduce methods to evaluate communication schemes by extracting meaningful symbols from messages, and measuring various properties of them. Toward establishing a typology of possible compositional communication schemes, we then establish reference points for emergent communication schemes based on both idealized artificial languages and natural languages.
4.1 Evaluating Communication Schemes
The need for stronger evaluations in EmCom is well-established Boldt and Mortensen (2024b); Chaabouni et al. (2020), and especially prominent in our problem setting, which aims to consider a greater number of plausible communication schemes than prior work. As such, we next introduce and appraise several algorithms for segmenting meaningful symbols from sequences of characters, as well as evaluation metrics for sequences of symbols in communication schemes.
4.1.1 Symbol Segmentation Algorithms
To extract meaningful, morpheme-like symbols from messages (essential for double articulation), we consider two approaches: one from prior work based on Harris’ articulation scheme Ueda et al. (2023), and a previously unexplored approach based on byte-pair encoding Gage (1994). We define the resulting segmented symbol vocabulary as the set union of symbols extracted from each message.
Harris’ Articulation Scheme
Harris’ articulation scheme (HAS) has been proposed as a desirable, naturalistic trait for emergent languages to exhibit Ueda et al. (2023); Ueda and Taniguchi (2024). HAS captures the tendency of word boundaries to be identifiable purely by character-level entropy. Specifically, the scheme claims that while entropy generally decreases across an utterance, it will increase at morpheme boundaries. This implies an algorithm for segmenting morphemes by drawing boundaries at indices where , where is a threshold that determines how sensitive the algorithm is. In all of our analyses, is set to 0 to maximize sensitivity. We use the implementation provided by Ueda et al. (2023).
Byte-Pair Encoding
HAS is a powerful tool for identifying symbols in communication schemes, but it is difficult to control the degree of segmentation by it. As such, we propose an alternative approach based on the Byte-Pair Encoding (BPE) algorithm Gage (1994), which has been widely used in tokenizers for NLP systems Sennrich et al. (2016). Importantly for our purposes, the properties of the subwords generated by BPE have been shown to enable characterization of languages by established typological categories, such as analytic and synthetic Gutierrez-Vasques et al. (2023). BPE is configurable with a maximum vocabulary size . Early compressions tend to be most informative Gutierrez-Vasques et al. (2021, 2023), so we apply BPE with a merging cutoff of 96, approximately twice the expected number of symbols (i.e., the total number of unique attribute values), and additionally apply BPE with maximum compression.
4.1.2 Evaluation Metrics
Given a segmented symbol vocabulary , we next introduce common metrics for its compositionality from prior work, along with methods to evaluate concatenativity and fusionality, more specific forms of compositionality that typical compositionality metrics do not directly measure.
Compositionality
Compositionality is the core objective for many EmCom works, motivated by interpretability and its prominence in natural languages. Topographic similarity (TopSim) is a common measure for this Brighton and Kirby (2006); Lazaridou et al. (2018), formally defined as the correlation of the distance (typically Levenshtein distance; Levenshtein, 1965) between pairs of messages and the distance between their meanings. This provides a global measure of how similar the symbols in messages for similar configurations.
However, TopSim places minimal restrictions on how similarity of messages is calculated. As such, Chaabouni et al. (2020) propose bag-of-symbols disentanglement (BoSDis), an alternative metric accounting for permutation-invariant communication schemes222They also propose positional disentanglement, which we omit but provide more details about in Appendix B. (e.g., in English, “red and blue” is equivalent to “blue and red”):
BoSDis calculates the average difference in mutual information between the count of a vocabulary symbol and attributes and , which respectively have the most and second-most mutual information with . This is normalized by entropy for . This metric captures the idea that a symbol’s presence in a message uniquely communicates a specific value of a specific attribute. This should generally not be true in our small-vocabulary setting, which necessitates reuse of characters in multi-character symbols to induce double articulation,333An obvious exception is one-character symbols, which are possible but can only comprise a small proportion of symbols in an accurate communication scheme. thus BoSDis should generally be low. However, if a symbol segmentation algorithm is effective at identifying symbols which uniquely refer to some attribute, the BoSDis with respect to those symbols should be higher. We thus use this as a gauge of whether HAS and BPE extract meaningful symbols through a ratio BoSDis of the BoSDis for a segmented symbol vocabulary to that for the unsegmented character vocabulary . If BoSDis, this indicates that creates a more disentangled representation than individual characters , suggesting successful identification of meaningful multi-character symbols.
Concatenativity
TopSim and BoSDis fail to account for non-trivial and variable forms of compositionality, including some present in natural language Korbak et al. (2020); Conklin and Smith (2023). For example, TopSim based on Levenshtein distance of characters reflects an assumption that characters should not be modified or reordered under composition. Contrarily, we aim to consider all schemes in which messages are formed via a systematic composition of symbols (i.e., as opposed to a memorized “hash function”) as compositional.
One particular form of compositionality we aim to tease apart from others is concatenativity, i.e., the degree to which composable symbols consist of uninterrupted streams of characters (analogous to phonemes) that are concatenated under composition. 444This property is closely related to the constraint of integrity in optimality theory, which prohibits output forms from containing multiple correspondents to an input form McCarthy and Prince (1995). Using English verb conjugation as an example, “played” is a concatenative message consisting of “play” and “ed.” By contrast, “sang” is a nonconcatenative message since the characters “s_ng” correspond to the root “sing,” but are interrupted by “a,” which communicates tense.
As HAS and BPE require symbols to be composed of consecutive characters, we estimate concatenativity as the average number of symbols in segmented messages of a communication scheme.555For HAS, we use the number of boundaries drawn. Note that concatenative compositionality implies compressibility via algorithms like BPE, but not necessarily vice versa. Since HAS relies on entropy, symbol length based on it may appear artificially low in communication schemes that include some random or otherwise non-meaningful characters. We refer to these length metrics as HASLen and BPELen|V|. A smaller number of symbols indicates greater concatenativity, as more consecutive characters belong in self-contained units of meaning. Of course, not all compositional communication schemes are concatenative. In our forthcoming experiments, we will apply concatenativity metrics to contrived concatenative and nonconcatenative compositional schemes (introduced in Section 5.1) to analyze how these properties relate.
Fusionality
Fusionality or degree of fusion is the expression of multiple attributes of meaning within one morpheme or symbol. Prior work examines this by taking conditional probabilities of surface forms after ablating feature pairs from the dataset Rathi et al. (2021, 2022). However, this requires training a model for each feature pair, a computationally expensive approach. The fusion of two attributes is equivalent to treating them as one attribute , for which symbols may be chosen arbitrarily, and a language may be compositional with respect to a particular fused attribute set. As such, we define a novel metric of fused TopSim (F-TopSim), calculated by the maximum TopSim over the fusions of all attribute pairs. A language is fusional if , which we summarize by a F-TopSim .
5 Artificial and Natural Reference Points
Next, we curate reference compositional communication schemes. To evaluate the ability of metrics to capture compositionality, concatenativity, and fusionality and establish signatures as expected ranges of metric, we first propose several contrived artificial languages to target these properties. To contextualize EmCom results with real-world communication schemes, we develop additional languages based on natural language inflections.
5.1 Artificial Languages
As shown in Figure 2, we generate five artificial languages using identical message length and character vocabulary size to Attr-Val settings. The first four types compose symbols with various levels of concatenativity, considering symbols of constant and mixed length. Notably, these languages preserve symbols and the ordering of their characters under composition, favored by TopSim. The fusion language, however, violates this by assigning symbols to combinations of attribute values rather than individual ones. These languages illustrate a range of morphological complexity.666Additional artificial languages explored in Appendix C. In Table 1, we report metrics for the compositionality (including effectiveness of symbol extraction), fusionality, and learnability of artificial languages. In Table 2, we report concatenativity metrics. We provide an extended discussion and additional artificial languages in Appendix C, summarizing key observations below.
| Language | TopSim | BoSDis | HAS BoSDis | BPE96 BoSDis | BPE BoSDis | F-TopSim | Epochs |
| Perf. Conc. | .628 | .076 | 4.889 | 7.836 | .751.34 | -.236.01 | 4.875 |
| Mixed Conc. | .621 | .076 | 2.146 | 1.829 | .430.23 | -.236.01 | 3.875 |
| Nonconcat. | .624 | .076 | .509 | .825 | .379.15 | -.237.01 | 4.875 |
| Var. Length | .451 | .085 | 1.965 | 5.110 | .796.28 | -.131.02 | 10.00 |
| Fusion | .238 | .126 | 9.685 | 4.474 | .402.03 | .175.01 | 29.13 |
| Random | .000 | .001 | 3.437 | 8.263 | 19.517.32 | .000.00 | 100+ |
| Language | HASLen | BPELen96 | BPELen |
| Perf. Conc. | 2.758 | 3.185 | 2.001 |
| Mixed Conc. | 3.791 | 4.489 | 2.178 |
| Nonconcat. | 4.558 | 4.722.07 | 2.294 |
| Var. Length | 2.463 | 3.078.18 | 1.995 |
| Fusion | 3.502 | 4.341.02 | 2.004 |
| Random | 1.593 | 4.997 | 3.105 |
First, the degree of concatenation in artificial languages has minimal effect on learnability, implying a lack of inductive bias toward concatenation in recurrent networks. BoSDis metrics indicate that HAS and BPE extract more meaningful morphemes from more concatenative languages, despite all artificial languages being compositional. HASLen and BPELen are lower in more concatenative languages, suggesting our proposed metrics effectively reflect concatenativity. Mixed concatenative languages have higher variance in BPELen, falling along a spectrum between perfect concatenation and nonconcatenation, and random languages have markedly higher asymptotic BPELen than systematic languages, further evidence of BPE’s effectiveness. However, as expected, HAS is less effective for languages with randomness, evidenced by mixed concatenative and variable length languages having relatively low HASLen and HAS BoSDis, and random languages having low HASLen.777Supplementary results with HASLen in Appendix D. BoSDis metrics themselves are consistently greater than 1 for random languages, suggesting an over-sensitivity to noise in the absence of other signal. Lastly, F-TopSim effectively reflects the fusionality of the fusion language.
5.2 Natural Languages
As a comparison to humanlike communication schemes, we also analyze subsets of the verb conjugations of two natural languages, Spanish and Arabic, as Attr-Val languages. We chose these due to the availability of high-quality datasets, the fact that both conjugate for tense and person, and the fact that they represent fairly divergent communication schemes along the concatenativity axis. Spanish data comes from the Jehle Spanish Verb Database Jehle (2011), and Arabic data was sourced from Nawaz et al. (2025). We select a set of verb attributes comparable to the inflection Attr-Val setting: root, tense, and person. We then generate Attr-Val configurations based on combinations of these verb attributes sampled from the data, which serve as cognitively plausible (albeit simplified) representations of actual cognitive objects and events encoded and reconstructed in natural communication. For other attributes beyond tense and person, we choose a constant value (e.g., mood is always indicative). We generate 50 such inflectional sublanguages for Spanish and Arabic, each of which are mappings between lexeme-slot tuples and surface forms , as per Wu et al. (2019).
We find that up to 76% of generated Spanish sublanguages can be meaningfully segmented by either HAS or BPE (i.e., BoSDis ratio greater than 1), while only up to 8% of the Arabic sublanguages can be.888Detailed analysis in Appendix E. The latter demonstrates how both tools for symbol segmentation are insufficient for nonconcatenative morphology.
6 EmCom Experimental Results
Our reimagined Attr-Val settings targeting double articulation and inflection, general and specific evaluation metrics for communication schemes, and illustrative artificial and natural language reference points enable us to ask previously unexplored research questions around morphology in EmCom. To begin this exploration, we specifically ask:
1.
Do more concatenative languages emerge under a human-inspired evolutionary pressure for ease of articulation?
2.
How do topographic similarity in emergent and natural languages for inflection compare?
3.
Do emergent languages exhibit fusionality in inflection similarly to natural languages?
In our experiments, agents are trained with eight different random initializations in the default and inflection Attr-Val games. Nearly all listeners exceed 90% accuracy on the task; more details are provided in Appendix F. Five of seven successful languages exhibit a BoSDis ratio greater than 1 under BPE96 in the default Attr-Val setting, while all do in the inflection setting. These results imply the general existence of multi-character meaningful segments in the emerged languages, confirming that double articulation is achieved.
6.1 Concatenativity and Ease of Articulation
Prior work has often explored pressures inspired by human evolution, such as ease of learning Kirby et al. (2014); Li and Bowling (2019) among others Vithanage et al. (2023); Galke and Raviv (2024), to encourage compositionality in emergent communication schemes. Meanwhile, prior work often assumes or favors concatenativity in their evaluations of schemes Resnick et al. (2020); Ueda et al. (2023) despite the possibility of compositional but nonconcatenative communication schemes (as demonstrated in our artificial and natural languages).
In order to achieve a goal of concatenative emergent languages with double articulation, we hypothesize that an appropriate pressure to explore may be ease of articulation. Human spoken languages are universally subject to phonological rules which drive the possibility of vocal sounds based on the environment and vocal tract physiology Dediu et al. (2017). The interface between phonology and morphology is a key area of study in linguistics, and a computational evolutionary model of morphology (as we hope to lay the groundwork for) would be incomplete without considering that morphemes are made up of meaningless components subject to similar rules independent of their meanings. Specifically, there are disallowed clusters of sounds in spoken languages (e.g., Hawaiian disallows consonant clusters), which may represent a significant pressure towards concatenative languages. To test this, we added a term to the training loss for a toy simulated phonology for agents, where even-valued characters can not occur next to even-valued characters, and odd-valued characters can not occur next to odd-valued characters.999More details provided in Appendix H.
Both without and with this pressure, we visualize the concatenativity (measured by BPELen) of emergent communication schemes in the default and inflection Attr-Val settings respectively in Figures 3 and 4. First, we observe no inductive bias toward concatenavity in our RNN-based agents as previous works implicitly assume; instead, emergent languages are distributed along a spectrum between perfectly concatenative and nonconcatenative. This may not be unrealistic: the natural languages also exhibit such a spectrum. Spanish appears relatively diverse, likely due to the presence of vowel patterns and auxilliary verbs. Arabic is more bimodal, with some very nonconcatenative sublanguages and some which have a high maximal compression due to being very fusional.
However, the introduction of articulation pressure sharply decreases BPELen in emergent languages. We perform additional -tests on the mean BPELen96 of emergent languages between conditions, observing statistical significance in this change for both the default and inflection Attr-Val setting ( and respectively). This suggests ease of articulation is indeed a strong pressure towards concatenation, and the presence of “unpronounceable” clusters at the phonological (character) level is sufficient to induce concatenation at the morphological (symbol) level. More broadly, this may help explain the dominance of concatenative morphology in natural languages.101010In Appendix J, we provide a thorough analysis of concatenativity and fusionality metrics with and without this pressure in both the default and inflection Attr-Val settings.
6.2 TopSim in Inflectional Languages
To shed more light on the compositionality of naturalistic communication schemes, we compare the TopSim of emergent and natural languages in the inflection Attr-Val setting in Figure 5, including emergent languages with the articulation pressure. Compared to natural languages, TopSim for emergent languages was low, except for one instance with a TopSim of .507 (in the range of natural languages). Interestingly, the articulation pressure resulted in a small decrease in TopSim, including one outlier language which had negative TopSim despite exceeding 90% reconstruction accuracy. Together, this suggests that emergent languages still have room for improvement in achieving humanlike topographic similarity, and there may remain more pressures toward this to identify and explore.
6.3 Fusionality Under Inflection
When reinterpreting Attr-Val as inflection, fusionality becomes a naturalistic trait for languages. Suppletion (fusion of a root with grammatical features, e.g., in went as the past tense form of go) is much rarer than fusion of two or more grammatical features. To judge whether this holds in EmCom, we identify the most fused attribute pair which exhibits the greatest F-TopSim in emergent and natural languages. As expected, attributes 1 and 2 (tense and person, which have the lowest number of values) had the highest F-TopSim in 6 of 8 emergent languages, suggesting these attributes are most likely to fuse. This mirrors the fusion of fixed grammatical features in natural language. In Arabic, most sublanguages (78%) also favored the fusion of these attributes. Unexpectedly, in Spanish, attributes 0 (root) and 1 (tense) fused most often (48% of sublanguages), with (1,2) at 30%. This may be due to the fact that specific roots can influence the form of inflectional affixes, e.g., the -amos suffix only applies to -ar verbs like llorar. An explanation for these trends may be that symbolic complexity Zhang (2025) increases the least in fusing these attributes, i.e., only 5 symbols are required to communicate the values for these two attributes before fusion, 6 are required after fusing them. Meanwhile, fusing the root attribute (with 42 values) with another attribute can drastically increase the symbolic complexity (i.e., 2-3 times).
7 Conclusion
In this work, we outline an experimental paradigm for emergent morphology, motivated by the principle of double articulation and a rich analogy between the Attr-Val game and inflectional morphology. We review existing metrics for EmCom, proposing a new metric for concatenativity based on BPE, and a new metric for fusionality based on variations in TopSims under fusion of attributes.
In our empirical analyses, we calibrate our metrics towards a wide array of idealized compositional schemes that are possible under this paradigm, as well as schemes based on natural language inflection. We then demonstrate through EmCom experiments the power of this reimagined setting to produce linguistically interesting analyses grounded in comparison to natural language. In particular, we show that the double articulation aspect of our paradigm reveals simulated articulatory constraints as a promising pressure toward concatenativity, which prior work has often implicitly assumed in evaluations of compositionality and attempted to elicit through other pressures, and future work may use to test more specific morphophonological hypotheses. Further, our emergent languages rederive the bias towards fusion of grammatical features in natural language. We hope that future work will adopt the small-vocabulary and inflectional Attr-Val settings for focused, linguistically motivated studies of morphology, and will benefit from being able to directly compare to natural language.
Limitations
Representativeness of natural languages.
In our natural language reference points, we only considered Arabic and Spanish, which are not a representative sample of the world’s natural languages, and were chosen in part due to the availability of data. Without a much larger and more diverse dataset, we cannot definitely say whether a given emergent language is naturalistic. Additionally, the Attr-Val scheme we focused on is a rather small subset of the complexity of natural morphologies. A full analysis of a large set of natural languages, to cover this diversity of morphologies, is outside the scope of this paper, but we believe our examples will provide a solid initial point of comparison for emergent languages. Boldt and Mortensen (2024a) explored larger-scale cross comparisons of emergent languages, and we anticipate that the addition of parallel natural language data to such analyses would prove useful for future work.
Representativeness of articulation pressure.
The choice of an even-odd toy phonological constraint for the articulation pressure was arbitrary, and not representative of the range of such constraints in natural language. An alternative explanation for the increased concatenativity from this pressure is that the predictable even-odd alternation pattern resulting from this pressure simply made the data more compressible due to the fewer number of possible variations in messages. While this is only a simple proof of concept of the sorts of linguistically motivated experiments we can do when embracing morphology and double articulation in EmCom, we argue that phonological constraints in natural languages may provide a similar advantage in human language learning, and encourage future investigation of how they may influence concatenativity and other properties in EmCom.
8 Acknowledgements
We thank our anonymous reviewers for their thoughtful and constructive feedback on this work in progress. We thank Zheyuan Zhang for helpful discussion about this work, and we thank Ryo Ueda for consulting with us on the implementation of metrics in Ueda et al. (2023).
Appendix A Harris’ Articulation Scheme
We replicated the results in Ueda et al. (2023) that C-TopSim is greater than W-Topsim. This result led the authors to suggest that Criterion 3 for Harris’ Articulation Scheme. (i.e. W-TopSim should be higher than C-TopSim) may be unsolvable.
| C-TopSim | W-TopSim |
| 0.611 | 0.362 |
As discussed in Section 4.1.2, we use ratios of BoSDis to judge whether segmented symbols are meaningful. We assert that this is a reasonable replacement for the intended purpose of the criterion, since it asks the question "do segments act as symbols which tend to refer to some meanings more than others."
Ueda et al. (2023) report negative results on the basis of low W-TopSim, which does not hold for our replacement metric of the BosDis ratio. Thus, we believe that emergent languages do exhibit HAS, and that HAS does not represent a significant gap between natural and emerged languages, as is claimed. The HAS algorithm, therefore represents a useful tool for interpreting emergent communication.
Appendix B Positional Disentanglement
Motivated by the possibility that symbols in different positions could indicate different attributes (common in natural languages), Chaabouni et al. (2020) propose positional disentanglement (PosDis):
For a communication scheme with messages of length , PosDis calculates the average difference in mutual information between the value of a symbol in position and attributes and , which respectively have the most and second-most mutual information with .111111Given the reliance on a consistent message length, we only apply PosDis based on individual characters rather than symbols extracted from HAS and BPE. This is then normalized by the entropy of . PosDis will be high if attributes are uniquely distinguishable by positions of symbols in messages. However, as we did not observe insightful trends with respect to this metric, we omitted PosDis from results presented in the paper to conserve space.
Appendix C Artificial Languages Extended Analysis
In this appendix, we provide extended discussions about concatenation and fusion in artificial languages. We then propose and evaluate additional nontrivial compositional artificial languages based on mutation and more general functional compositionality.
On Concatenation
First, the level of concatenation in the first three artificial languages has insignificant effect on TopSim and learnability, implying a lack of inductive bias toward concatenation in recurrent listener neural networks. Notably, this is not consistent with human results in linguistics Finley and Newport (2021), and implies additional pressures may be required to induce concatenative morphology. Meanwhile, as expected, BoSDis metrics indicate that both HAS and BPE successfully extract more meaningful morphemes from more concatenative languages, despite all artificial languages being compositional. In the mixed concatenative and variable length languages, HAS yields lower metrics, suggesting it may indeed be less effective in languages with randomness. BPE can seemingly compensate for this as long as a sufficiently limited vocabulary size is chosen. The right vocabulary size is not obvious, however, as the BosDis for BPE is relatively low for artificial languages, suggesting an over-segmented language that erases the disentanglement of symbols. While segmentation algorithms behave as expected for systematic languages, these metrics are greater than 1 for random languages. This may result from over-sensitivity to noise in the absence of other signal, e.g., the presence of a few rare, highly predictive symbols may skew the metric upwards. This may also be attributed to a vanishingly small BoSDis for character-level segmentation. Future research should consider this edge case when comparing BoSDis before and after symbol segmentation.
Focusing in on concatenativity as measured by HASLen and BPELen metrics, we unsurprisingly observe that the perfectly concatenative and variable length languages (which are both entirely concatenative) exhibit the highest concatenativity, except for BPELen, which shows no sensitivity to the level of concatenativity in artificial languages. This further supports the possibility of over-segmenting languages into symbols. As such, in Figure 6, we visualize the BPELen for a range of symbol vocabulary sizes. While minimal compression and maximal compression are similar across all artificial languages except the random language, there is a wide range in which BPELen behaves as expected. The mixed concatenative languages appear to fall on a spectrum between perfect concatenation and nonconcatenation, thus BPELen appears to be an effective measure of concatenativity. Meanwhile, due to the marked difference in BPELen between random and systematic languages, maximal compression may be an effective way to judge the amount of randomness in messages.
On Fusion
F-TopSim values indicate that the fusion language is indeed the most fusional. We observe that F-TopSim for random languages is zero, which does not suggest fusionality, rather that TopSim does not change with respect to fusion of attributes (expected for a random language, which should not assign any meaningful symbols to any attribute values or pairs thereof). We observe that fusionality does not vary significantly with respect to concatenativity, but increases in the variable length concatenative language.
On Mutation and Functional Compositionality
Traditional metrics for compositionality such as TopSim do not account for some types of nontrivial compositionality, such as phonemic mutation (e.g., of vowels in inflection of sing, sang, and sung), or more general function-based compositionality, where symbols from some attributes serve as functions applied to other attributes. As visualized in Figure 7, we defined additional languages for these phenomena. Specifically, following Korbak et al. (2020), the mutation language allows symbols to overlap, and combines overlapping characters’ values by addition modulo . We specifically consider two forms of the mutation language: Mutation3, which consists of length 5 symbols each overlapping each other at 3 character positions, and Mutation0, which consists of length 9 symbols each entirely overlapping. Lastly, the symbolic function language defines one symbol as a function, specifically a reordering function applied to the characters of another symbol. Beyond these artificial languages, we also consider a baseline random language.
On BoSDis in random languages
One may note that the BoSDis for random languages is quite high. We expect this for 2 reasons:
1.
They have character-level BoSDis close to 0.
2.
By sheer random chance, some segments of characters will be better predictors of attributes than others. This results in a slight increase in symbol-level BoSDis. This results in a very high ratio, since even a small absolute increase may be 10x as great as the original BosDis. For this reason we don’t advocate putting much stock in the magnitude of the ratio.
Whether this is a problem for the metric is a matter of interpretation. We believe that since random languages are very unique when all metrics, especially TopSim, are taken into account, the risk of misinterpreting a random language as a meaningfully segmented one is low. Furthermore, this enhances the confidence we can place in low BosDis ratio being an indicator of non-concatenativity, as these languages are qualitatively different from both random and concatenative languages (see BoSDiS< 1 for perfectly nonconcatenative languages in Table 1).
The evaluation metrics for these artificial languages are listed in Tables 4 and 5. Despite being compositional, these languages drastically impact evaluation metrics. The reordering and Mutation3 language are substantially less compositional (according to TopSim and BoSDis), segmentable, and learnable than fully concatenative languages. In Mutation0, where all characters are mutated twice, the languages were essentially indistinguishable from random languages by almost all metrics, including learnability. Even though this language consists entirely of messages formed by composing symbols according to simple, defined rules, in practice this compositionality may not matter or be recognized through existing metrics. The only metrics where there was a noticeable difference are the BoSDis ratios for HAS and BPE, but the interpretations of these are unclear, since segments cannot be disentangled symbols for this type of language.
| Language | TopSim | BoSDis | HAS BoSDis | BPE96 BoSDis | BPE BoSDis | F-TopSim | Epochs |
| Reordering | .227 | .042 | 23.24 | 88.64 | 13.44 11.84 | -0.0060.01 | 20.25 |
| Mutation3 | .359 | .060 | 40.37 | 17.31 | .578 | -.023.01 | 24.889.85 |
| Mutation0 | .000 | .001 | 3146 | 7461 | 17.2045.46 | .001.00 | 100+ |
| Language | HASLen | BPELen96 | BPELen |
| Reordering | 2.268 | 4.127 | 2.086 |
| Mutation3 | 4.843 | 4.461 | 2.273 |
| Mutation0 | 1.560 | 4.998 | 3.107 |
Appendix D HASLen for Natural and Emergent Languages
In Figure 8, we visualize the HASLen for emergent languages (on average) along with Spanish and Arabic natural language reference points. Following observations in Section 5.1, we observe that Spanish and Arabic are scored with a higher HASLen than random languages. This further demonstrates HAS’ sensitivity to randomness, making it difficult to use for reference points of concatenativity (despite its effectiveness in identifying symbol boundaries).
Appendix E Segmentation Analysis of Natural Languages
As shown in Table 6, when calculating the BoSDis ratio for these languages, we find that up to 76% of generated Spanish sublanguages121212Some failures here could be attributed to nonconcatenativity in the dataset caused by prepended auxiliary verbs. can be meaningfully segmented by either HAS or BPE (i.e., ratio greater than 1), while virtually none of the Arabic sublanguages can be. On the latter, this demonstrates how existing tools for symbol segmentation are insufficient to capture highly nonconcatenative morphologies.
| HAS | BPE96 | |
| Spanish | 0.76 | 0.74 |
| Arabic | 0.08 | 0.00 |
Appendix F Agent Implementation Details and Task Accuracy
In line with prior related works, we implement the sender and receiver agents as single layer gated recurrent unit (GRU; Cho et al., 2014) models using the EGG framework Kharitonov et al. (2019), with hidden dimensions of 500. Observed states were implemented by concatenating one-hot vectors for each attribute. Senders are optimized using REINFORCE Williams (1992) and receivers with backpropagation following Chaabouni et al. (2019) using the accuracy (i.e., proportion of attributes correctly reconstructed) as a loss, with cross-entropy loss as a supplementary loss term. It was found in initial experiments that this allowed smoother and more reliable convergence with lower sequence lengths, which is desirable for obtaining interpretable and minimally complex languages. Each run occurred on a single A40 GPU with 48GB VRAM and took about 8 hours, although many models converged before then. Models generally converged around the 400 epoch mark.
In the inflection Attr-Val setting, we additionally weighted the root accuracy so it represented 90% of the accuracy optimized by the inflection-condition agents, for two reasons: initial experiments revealed unstable training without weighting, and this reflects that far more of the semantic content of a word is stored in the lexeme, rather than the slot, so the lexeme is more important to communicate. We also experimented with larger configurations with more “roots,” but we found that higher input sizes led to unstable training, where agents would learn either to inflect or distinguish roots, but not both. We attribute this to the increased symbolic complexity of the setting Zhang (2025).
We ran experiments with agents 8 times for each setting with different random initializations. In the default Attr-Val setting, 7 of 8 emerged languages were successful, i.e., achieved greater than 90% accuracy on the reconstruction task. For other experiments, including with the articulation pressure discussed in Section 6.1 and the inflectional Attr-Val setting, all languages were considered successful.
Appendix G Ease-of-Learning Pressure Experiment
Inspired by prior work in iterated learning Kirby et al. (2014), we implement an experiment incentivizing ease of learning in emergent languages Li and Bowling (2019). Specifically, every 100 epochs, we reset the listener’s network weights. In theory, this should encourage emergent languages that are easier for the listener to learn quickly, which Li and Bowling show improves compositionality of the resulting languages. To understand whether this ease of learning is attributed to concatenativity, which prior work has typically not distinguished from compositionality, Figure 9 visualizes the BPELen of emergent languages with this pressure in the default Attr-Val setting (i.e., 3 attributes each with 16 values). Interestingly, we find that the languages resulting from this pressure are not significantly more concatenative than those without it shown in Figure 4. Such a result is unexpected from an intuitive standpoint, but aligns with the results in 6, which show that concatenativity, by itself, does not have a relationship with learnability. This motivates future research around how various pressures may influence fine-grained properties of emergent languages, and how much of human language structure can be explained by the generational (iterated) learning paradigm.
Appendix H Articulation Pressure Details
Formally, for a message comprised of a sequence of integer tokens, the loss term for ease of articulation pressure is computed as:
is a hyperparameter which controls how ’strict’ the phonological rule is. We tried 1, 10, and 100 as values and found that 10 resulted in fairly strict adherence without incurring optimization problems, so in our reported experiment.
8 of 8 seeds in the default setting succeeded, and 7 of 8 of the inflectional setting. In the default Attr-Val setting, all of the successful articulatory languages exhibited . In the inflection setting, 3 of 7 of the successful languages exhibited , indicating that not all of the segmentations were meaningful. We hypothesize that the additional constraint caused difficulties in segmenting, or alternatively that the pressure may have also resulted in highly nonconcatenative languages that appear concatenative due to the additional pattern. The fact that the intersection of these two conditions results in a somewhat unintuitive result underscores the need for future work to understand the fine-grained details of EmCom.
Appendix I Additional Notes On Inflection Setting
In the inflection setting of Attr-Val, one possible limitation is the lack of control on the number of combinations of attribute values (rather than unique number of attribute values, which we control for instead). We conducted a preliminary experiment where we controlled for number of combinations rather than number of values, however these models universally failed to converge, instead learning to distinguish either grammatical features or roots, but never both.
We believe that this failure to converge is related to the model capacity, as explored by Resnick et al. (2020), Specifically, we hypothesize that for languages with more unique values (i.e., symbolic complexity as defined in Zhang, 2025), the minimum model capacity for a successful language increases. While we could simply increase the size of this model (e.g., embedding and/or hidden size) to better encourage its convergence, it would no longer be comparable to other results. Meanwhile, if we increased the sizes of all models, this may influence compositionality in less complex configurations. As such, we decided that a comparison controlling for the number of combinations would not be possible within the scope of this paper.
Nevertheless, we chose our specific Attr-Val configuration in the interest of comparability to natural language. To make a good comparison, we needed to be able to select features from natural language with equal or greater cardinalities to our artificial features. As such, there are only so many grammatical features we can introduce and find high-quality data for, and considering features with larger cardinalities would preclude a lot of natural language features like gender, number, person, etc. and severely limit our dataset. The specific values were chosen to allow this comparison and control for symbolic complexity.
| Setting | Articulation? | BPELen_96 | F-TopSim | HAS BoSDis | BPE BoSDis |
| 16x16x16 | Y | 3.534 | -0.070 | 186 | 12.1 |
| N | 3.825 | 0.0112 | 3.25 | 1.25 | |
| 42x2x3 | Y | 2.340 | -0.0083 | 1.56 | 30.7 |
| N | 3.257 | -0.0078 | 113 | 45.4 |
Appendix J Concatenativity and Fusionality Metrics for all EmCom Experiments
In Table 7, we summarize the concatenativity and ease of fusion metrics for each combination of Attr-Val setting and presence of the articulation pressure. As observed there, the average F-TopSim values are universally close to 0 in emergent languages, indicating some inconsistent degree of fusion. Only in the default setting without articulation pressure is F-TopSim positive on average, however, indicating a high degree of fusion. A t-test for mean F-TopSim between articulation pressure and no articulation pressure in the default setting yields highly significant results (p << 0.05). A possible explanation for this is that fusion creates more total symbols, and articulation pressure discourages extra symbols, as there are only so many readily distinguishable symbols that abide by the articulation rule.